OpenStudy (anonymous):

find y' by implicit differentiation a. 4x^2+9y^2=36 b. .5x+.5y-4=0

myininaya (myininaya):

\[(4x^2)'=?\]

myininaya (myininaya):

What is derivative of 4x^2?

OpenStudy (anonymous):

8x

myininaya (myininaya):

(9y^2)'=?

OpenStudy (anonymous):

18x

myininaya (myininaya):

18y*y' <----because the derivative of y is y'

myininaya (myininaya):

So we have \[8x+18yy'=0\] Solve for y'